# Atlas: A Framework for ML Lifecycle Provenance & Transparency

### v1

@@ Abstract
The rapid adoption of open source machine learning (ML) datasets and models exposes today’s AI applications to critical risks like data poisoning and supply chain attacks across the ML lifecycle. With growing regulatory pressure to address these issues through greater transparency, ML model vendors face challenges balancing these requirements against confidentiality for data and intellectual property needs.
We propose Atlas, a framework that enables fully attestable ML pipelines. Atlas leverages open specifications for data and software supply chain provenance to collect verifiable records of model artifact authenticity and end-to-end lineage metadata. Atlas combines trusted hardware and transparency logs to enhance metadata integrity, preserve data confidentiality, and limit unauthorized access during ML pipeline operations, from training through deployment.
Our prototype implementation of Atlas integrates several open-source tools to build an ML lifecycle transparency system, and assess the practicality of Atlas through two case study ML pipelines.
@@ 1 Introduction
In recent years, machine learning (ML) models, have become increasingly popular. The pervasive use of large language models (LLMs), in particular, and multi-stakeholder involvement in model creation and deployment exacerbate security and privacy risks. These considerations are emphasized by the global nature and the complexity of large-scale ML deployments with different lifecycle stages:
  1. 1)
    Collection and sanitation of a training dataset from several public and proprietary sources.
  2. 2)
    Provisioning of the training environment (hardware and software).
  3. 3)
    Execution of training across many data centers.
  4. 4)
    Construction of a testing dataset from several sources, and the evaluation.
  5. 5)
    Deployment and use of the model for inference that is compliant with local laws or corporate policies.
Each of these stages is vulnerable to malicious or dishonest parties. For example, data can be poisoned [1, 2] during collection or training. Service providers executing outsourced training can shorten or omit critical steps to reduce their cost. Model providers can serve smaller models in SaaS, or even distribute malicious ones.
On the other hand, responsible model builders and other stakeholders may be incentivised or required to provide security and trust guarantees. They may want to prove low bias in their training data, offer easily verifiable performance claims, or guarantee end-to-end integrity of the model creation in high risk domains.
To address these challenges, it is necessary to guarantee the integrity of the entire ML lifecycle – beginning with the data, through the training, and finally, the evaluation and deployment. Was the data modified? Did the hardware and software environment adhere to the specification? Did the contractor follow the specified training procedure? Can I trust the evaluation? How can I guarantee that I am interacting with the intended model? These are example questions that showcase the breadth of the involved challenges that must be tackled to provide end-to-end security.
In this work, we introduce Atlas, a framework for enhancing the security and transparency of the lifecycle of ML models. Atlas establishes the baseline of fundamental components and capabilities needed for comprehensive provenance tracking at each stage of the ML lifecycle. Subsequently, Atlas defines the core integrity requirements for verifiable ML lifecycle transparency. We provide a reference implementation that instantiates Atlas using hardware-based security mechanisms – with trusted execution environment (TEE), including attestations.
@@
    We instantiate Atlas using TEEs and metadata-based provenance tracking.
@@
    We evaluate our Atlas prototype through two case studies: 1) fine-tuning of a BERT model [3, 4]; 2) fine-tuning of a bge-reranker model [5] .
@@ 2 Background & Related Work
2.1 Data Provenance & Authenticity
While not a new concern, the authenticity, provenance and attribution of media has received renewed attention in recent years due to the wide online proliferation of images, videos and text that have been manipulated or forged using generative ML models [6, 7].
One common approach for addressing digital image and video manipulation involves cryptographic techniques such as hashing, digital signatures and blockchains, which enable provenance authentication, tamper detection, and checking for divergent views of the content [8].
@@ 2.1 Data Provenance & Authenticity
Most notably, the Coalition for Content Provenance and Authenticity (C2PA) specification [9, 10] introduced by the Content Authenticity Initiative (CAI) [11] and Project Origin [12] provides cryptographic binding of authenticity claims, verifiable assertions about content origin and modifications, and tamper-evident audit trails [13].
Initially launched as a response to the growing challenge of deepfakes  [11] and digital content manipulation, the C2PA standard first gained traction in digital photography and journalism workflows. The standard’s extensible and interoperable metadata format has made C2PA an attractive option for ML dataset and model provenance tracking [14].
Other prior work in this space builds upon digitally signed metadata by leveraging distributed ledger technologies to create transparent and immutable content or provenance records. Examples of this approach include AMP [7], the News Provenance Project [15] and Collomosse et al. [14].
Atlas borrows techniques from C2PA and similar approaches, but integrates them directly into the ML system to track model artifact provenance closer to where the transformations occur to achieve higher metadata fidelity, rather than applying them in an ad hoc fashion after a particular dataset or model artifact has been created.
The Open Source Security Foundation (OpenSSF) Model Signing project [16] is a parallel effort in this space focusing specifically on cryptographic signing of model artifacts to ensure their authenticity and integrity. As such, it serves as a crucial building block but requires integration with other tools for more comprehensive supply chain security. While Model Signing might integrate with Atlas, additional mechanisms would be needed to address our goal of tracking full model lineage.
LakeFS [17] combines Git-like semantics with concepts from object stores such as S3 to provide a version control system for data, including ML datsets. Thus, LakeFS aims to capture data lineage by tracking changes to stored data over time, and allowing ML applications to reference specific versions of the stored data. This approach is meant to integrate with existing first-party data processing pipelines, but does not facilitate verification of data provenance by downstream consumers. Atlas’ metadata centered approach, on the other hand, enables first- and third-party ML dataset consumers to track changes and check their provenance, even when they may not have direct access to the data.
2.2 Supply Chain Integrity
While the ML supply chain is an emerging area of research, software supply chain security and integrity has been the subject of much prior work. Thus, several of these approaches are being explored for their applicability to enhancing the integrity of the ML model lifecycle.
Bills of Materials (BOM) have been employed to document the list of components of a hardware or software product for over three decades [18]. Due to recent regulations [19, 20], Software BOM (SBOM) have been the focus of many industry and academic efforts seeking to facilitate tracking software dependencies and other metadata to detect supply chain threats [21].
The AIBOM framework [22, 23] incorporates concepts from traditional software supply chain management while focusing on intended ML model use cases. Like SBOM, AIBOM provide a mechanism for tracking model software dependencies and maintaining model metadata. At the time of writing, we are not aware of any frameworks other than Atlas that utilize any sort of BOM data format to track ML model components.
Complementing BOM, a number of efforts like OpenSSF Supply Chain Levels for Software Artifacts (SLSA) [24] and SPDX Build [25] have sought to collect build provenance, i.e., metadata describing the process through which a particular artifact was produced. With its incremental guidance on implementing integrity measures for software builds and its build provenance data format, the SLSA framework [24] is gaining traction among software vendors, and also beginning to see some use for ML model fine tuning [26]. Atlas’ configurable metadata collection capabilities would enable us to collect build provenance at any stage of the ML model lifecycle, providing a more comprehensive view.
A number of frameworks for capturing and verifying a variety of security claims and metadata about the supply chain have been proposed. in-toto [27] collects authenticated claims across supply chain steps, including SBOM and SLSA metadata. In particular, in-toto enables software development pipeline owners and downstream artifact consumers to specify end-to-end supply chain policies, and validate that only the expected parties carried out specific steps in the supply chain, and artifacts underwent transformations in the expected order. Given recent and upcoming enhancements that further generalize the framework, in-toto may be a suitable option for specifying and verifying end-to-end ML model lifecycle integrity policies in Atlas.
Sigstore [28] provides a transparency log-based infrastructure for issuing signing credentials and validating digital signatures on supply chain artifacts and metadata. Similarly, Supply Chain Integrity, Transparency and Trust (SCITT) [29] is an architecture for implementing distributed ledger to provide global visibility and auditing for supply chain operations and claims. The SCITT architecture also includes confidential computing technologies to help ensure that only authorized parties submit claims to the transparency ledger.
Atlas builds upon several of these supply chain integrity and transparency approaches, but seeks to support multiple types of artifact provenance, supply chain metadata and integrity verification mechanisms all within a customizable, integrated framework designed for the ML model lifecycle.
2.3 Model Lineage Tracking
@@
The EQTY Lineage Explorer [30] provides granular tracking of model artifacts throughout the training process. The tool captures relationships between datasets, model checkpoints and hyperparameters, enabling developers to trace model evolution. Unlike Atlas, it lacks cryptographic guarantees for artifact authenticity and focuses primarily on manually collected development-time lineage, rather than automatically capturing and linking information across the entire ML lifecycle.
@@
ML experiment trackers like Weights and Biases [31], Neptune [32] and Kubeflow Pipelines [33] offer detailed run-time logging of model metadata about training runs, metrics, and model artifacts. While comprehensive for collaborative ML application development, these tools do not necessarily integrate transparently with common ML frameworks, and they typically provide only unauthenticated metadata that makes it less appropriate for establishing trust in a particular ML pipeline. Atlas, on the other hand, focuses making model lineage verifiable and on enabling adoption by supporting integration into existing open ML frameworks.
2.4 Hardware-Based Security for ML
@@
Recent developments in trusted execution environments (TEEs) technologies have made it more practical to run large-scale systems and workloads. Thus, a growing number of proposals leverage TEEs to harden a variety of ML systems through hardware-enforced isolation and integrity verification mechanisms.
@@
Chrapek et al. [34] use TEEs in a large language model (LLM) fine-tuning system running inside secure enclaves that help protect LLM code and data while in use. To this end, they evaluate their LLM system using the process-level Intel®Software Guard eXtensions (Intel®SGX) and virtual machine-level Intel®Trust Domain eXtensions (Intel®TDX) TEEs, showing that they are able to maintain practical performance overhead of less than 10% for models like Llama2 in both TEE configurations.
Laminator [35] demonstrates the application of TEEs to ML property attestation for ML model regulatory compliance verification. Specifically, this work leverages TEEs’ hardware-assisted remote attestation and cryptographic binding capabilities to generate verifiable ML model property cards. This work is complementary to Atlas and may enable us to extend our framework to support such ML property attestations.
@@
Mo et al.’s survey [36] evaluates 38 works that use various TEE implementations to enhance the privacy and integrity of ML training and inference operations. The survey highlights several gaps that existing TEE-based frameworks do not address, including the protection of a full ML lifecycle, which is the primary focus of Atlas.
@@ 3.1 Terminology
In Atlas, an ML model is composed of a number of artifacts that include the training dataset, ML algorithm, ML framework (e.g., PyTorch), model configuration (e.g., hyperparameters, weights), and metadata (e.g., license).
@@
The ML pipeline defines the sequence of operations or a workflow that transforms or operates on a model artifact at a particular stage of the ML lifecycle [37]. To support the standardization and repeatability of a specific workflow, the pipeline is also used to facilitate workflow management and automation, e.g., training or evaluation pipelines.
@@ 3.2.1 Stakeholders
The Artifact Producers are individuals and organizations that create ML model artifacts to provide or sell them to other parties. Since the artifacts may represent intellectual property and/or make use of personally identifying information (PII), producers have business and regulatory reasons to preserve the confidentiality of their artifacts. To save costs, artifact producers often outsource the operation of ML systems to external MLaaS providers.
ML-as-a-Service (MLaaS) providers operate and maintain the compute infrastructure needed to run ML systems on behalf of artifact producers. MLaaS providers may offer ML-specific services that leverage general-purpose systems (e.g., [38, 39, 40]), or provide special-purpose systems like orchestrators [33] or CI/CD (e.g. [41]) that can be used to build and run third-party ML systems.
A Hub is a system that stores and distributes model artifacts. Thus, model pipelines typically ingest and output artifacts to and from hubs during their execution, enabled by interfaces exposed by MLaaS providers.
Hubs may be operated by artifact producers themselves or by third-party vendors, containing open or closed source artifacts (e.g.,HuggingFace [42], PyTorch Hub [43]).
A Transparency Service in Atlas is responsible for generating, storing and distributing the information necessary to verify the authenticity and integrity of model artifacts, ML systems and lifecycles as a whole. We envision model vendors and independent parties to operate transparency services in practice.
Verifiers are entities that seek to evaluate or audit a particular ML model’s lifecycle with the goal of checking whether the model it produced was tampered with in an unintended or malicious way. To this end, verifiers in Atlas define a set of expectations about a producer’s model artifacts, MLaaS providers, and their ML system components. In practice, model users, vendors or regulatory entities may act as verifiers.
@@ 3.2.2 ML Lifecycle
In Atlas, we consider four common high-level stages in the ML lifecycle. Each builds upon the outputs and feedback from the others, forming a continuous cycle that allows models to evolve and adapt based on real-world usage.
1. Data processing: Raw data is collected, curated, processed into smaller units (e.g., tokens) and collated into a structure ingestable during training.
2. Training: a training algorithm processes a given dataset using an ML framework. The output of the training pipeline is an ML model that may undergo further fine tuning and evaluation, prior to being deployed for inferencing.
3. Evaluation: following training, model properties like its performance and accuracy are checked and further adjusted, prior to production release and deployment.
4. Deployment: Once an ML model has been trained and evaluated, it is deployed to a production system configured for inferencing. The new data obtained from clients during inference is sent back to a data processing pipeline to enhance the training dataset, and the model.
@@ 3.3 Threat Model
Atlas is concerned with addressing risks to ML model artifacts introduced via the ML lifecycle, i.e., attack vectors in third-party MLaaS and hubs across different ML pipelines that lead to compromised ML artifacts. Specifically, while transparency services, verifiers and model users are trusted in Atlas, we consider threats by MLaaS providers, hubs and artifact producers.
The adversary’s goal is to produce a tampered artifact or model, e.g., containing a hidden malicious component, so that a trusted transparency service generates a legitimate signature on the artifact or its metadata.
Compromised MLaaS providers and hubs may involve malicious insiders, or external adversaries seeking to subvert these systems by exploiting vulnerable components. Given their central position, MLaaS providers and hubs may thus be able to compromise model integrity at various stages of the ML lifecycle.
For example, a malicous MLaaS provider can poison the training data during the curation step of the data processing stage leading to backdoors [2, 1]. A malicious hub may, for instance, present a dataset or model with a mismatched signature (e.g., to a different model, or any of its component artifacts) to a model user or MLaaS provider, so that pipelines in subsequent stages of the ML lifecycle may ingest compromised dependencies.
As a result, these compromises may propagate through the ML lifecycle if they go undetected, ultimately leading to vulnerable ML models at the deployment stage. This risk is exacerbated if a hub colludes with an MLaaS providers to introduce or accept compromised ML pipeline inputs.
Artifact producers, on the other hand, may seek to compromise ML models in order to bypass regulatory requirements, introduce exploitable vulnerabilities or steal private information for profit (e.g, [44]). Thus, producers may collude with other untrusted participants, or inaccurately declare their dependencies leading to intentional omissions or misrepresentations their artifacts, to surreptitiously undermine the integrity of their artifacts.

### v2

@@ Abstract
We propose Atlas, a framework that enables fully attestable ML pipelines to address ML supply chain risks. Atlas leverages runtime pipeline monitoring and open specifications for data and software provenance to collect model artifact integrity and end-to-end lineage metadata. Atlas combines trusted hardware and transparency logs to enhance metadata integrity and enable efficient verification of ML pipeline operations, from training through deployment. Our prototype implementation of Atlas integrates open-source tools to build an ML lifecycle transparency framework.
@@ 1 Introduction
In recent years, machine learning (ML) models, have become increasingly popular. The pervasive use of large language models (LLMs), in particular, and multi-stakeholder involvement in model creation and deployment exacerbate security and privacy risks. These considerations are emphasized by the global nature and the complexity of large-scale ML deployments with different lifecycle stages (e.g., training dataset collection, execution of training). Each stage is vulnerable to malicious or dishonest parties. For example, data can be poisoned [1, 2] during collection or training. Service providers executing outsourced training can shorten or omit critical steps to reduce their cost. Popular model hubs hosting pre-trained models are vulnerable to compromises that may result in corrupted, reduced, or malicious model distributions [3, 4]
On the other hand, recent regulations [5, 6] require model builders and other stakeholders to provide evidence of ML model security and trust. They may need to prove low bias in their training data, offer easily verifiable performance claims, or demonstrate end-to-end integrity of model creation in high risk domains.
To address these challenges, integrity of the entire ML lifecycle must be recorded verifiably – beginning with the data, through the training, and finally, the evaluation and deployment. Was the data modified? Did the hardware and software environment adhere to the specification? Did the contractor follow the specified training procedure? Can I trust the evaluation? How can I guarantee that I am interacting with the intended model? These are example questions that showcase the breadth of the involved challenges that must be tackled to provide end-to-end security.
We introduce Atlas, a framework for enhancing the security and transparency of the lifecycle of ML models. Atlas establishes the baseline of fundamental components and capabilities needed for comprehensive provenance tracking at each stage of the ML lifecycle. Thus, rather than preventing attacks entirely, Atlas detects tampering by verifying the ML lifecycle.
Atlas addresses two challenges unique to ML lifecycle transparency. First, in contrast to the clear dependency trees of traditional software, datasets and algorithm code are tightly coupled in ML models, creating significantly more intricate provenance graphs [3]. Attesting the integrity of these relationships requires mechanisms that can track cross-organizational transformations across heterogeneous artifacts of varying sizes and formats [7].
Hence, Atlas monitors ML pipelines during execution and automatically collects ML system information using several data and software provenance frameworks. To strengthen the integrity of provenance generation, Atlas relies on hardware trusted execution environments.
Second, models are not static binaries – they can be further customized by downstream users. That is, ML models undergo a non-linear lifecycle where deployment results often necessitate refinement, creating feedback loops between inferencing and data processing [8].
This iterative process requires special adaptations to provenance tracking mechanisms. Atlas represents these non-linear development paths and enables cryptographic auditing using Merkle trees [9].
@@
    We instantiate Atlas using Intel Trust Domain eXtensions [10] and metadata-based provenance tracking.
@@
    We evaluate our Atlas prototype through a fine-tuning case study with a BERT model [11, 12].
@@ 2 Background & Related Work
Data Provenance & Authenticity. Provenance and attribution of media has recently received attention due to the online proliferation of manipulated or forged content using generative ML models [13, 14]. Prior work relies primarily on cryptographic hashing and digital signatures to provide data authenticity and integrity.
@@
The Coalition for Content Provenance and Authenticity (C2PA) specification [15, 16] digitally signs assertions about content origin to provide tamper-evident data audit trails [17, 18].
C2PA’s extensible metadata format also makes it suitable for ML model artifacts [19].
The Open Source Security Foundation (OpenSSF) Model Signing project [20] is a parallel effort focusing on integrity and authenticity of trained models. Other prior work [14, 21, 22] builds upon hashing and signing with distributed ledger technologies to create transparent and immutable content or provenance records.
These techniques are crucial building blocks for verifying ML model artifact authenticity and provenance, but each alone is insufficient for end-to-end ML supply chain transparency. In contrast, Atlas aims to integrate such techniques into ML systems to track model artifact provenance directly where the transformations occur.
Supply Chain Integrity. Recent cybersecurity regulations [23, 24] have shifted industry focus toward detecting supply chain threats via software dependency tracking with SBOMs [25]. Similarly, the AIBOM framework [26, 27] focuses on ML model supply chain management.
Complementing BOM, efforts like OpenSSF Supply Chain Levels for Software Artifacts (SLSA) [28] and SPDX Build [29] collect build provenance, i.e., metadata describing how a particular artifact was produced.
This approach is also being considered for ML model fine tuning [30]. Building on such supply chain metadata efforts, a number of frameworks [31, 32, 33] provide mechanisms for collecting, digitally signing and verifying authenticated claims across supply chain steps.
Atlas borrows concepts from supply chain integrity to support multiple types of software artifact provenance at any stage of the ML lifecyle, providing a more comprehensive view of a model’s supply chain. Works about evidence for other properties of the ML lifecycle such as assurance [8] are complementary.
Model Lineage Tracking.
@@
The EQTY Lineage Explorer [34] tracks model artifacts throughout the training process, capturing relationships between datasets, model checkpoints and hyperparameters. However, unlike Atlas, it lacks cryptographic authenticity properties and focuses primarily on manually collected development-time lineage, rather than automatically capturing and linking information across the entire ML lifecycle.
@@
ML experiment trackers like Weights and Biases [35], Neptune [36] and Kubeflow Pipelines [37] offer detailed run-time logging of model metadata about training runs, metrics, and model artifacts. These tools do not integrate transparently with common ML frameworks, and they typically provide only unauthenticated metadata. Atlas, on the other hand, seeks to make model lineage verifiable and support integration into ML frameworks like PyTorch [38].
@@
Hardware-Based Security for ML. Recent developments in trusted execution environment (TEEs) technologies have made it more practical to run large-scale systems and workloads [39, 40], including ML pipelines.
@@
Chrapek et al. [41] deployed and optimized a large language model (LLM) inside a TEE, showing how secure enclaves help protect LLM code and data while in use. They maintain practical performance in two TEE configurations based on Intel Software Guard eXtensions (Intel SGX) [42] and Intel Trust Domain eXtensions (Intel TDX) [10].
Laminator [43] and PraaS [44] demonstrate the application of TEEs to ML model or dataset property attestation and verification. Several efforts [7, 45, 46] use TEEs to build confidentiality frameworks for different ML lifecycle stages. These works are complementary to Atlas and may enable us to extend our framework.
@@
Mo et al.’s survey [47] evaluates 38 works that use various TEE implementations to enhance the privacy and integrity of ML training and inference operations. The survey highlights several gaps, including the protection of full ML lifecyles, which is the primary focus of Atlas.
We provide additional background in Appendix A.
@@ 3.1 Terminology
In Atlas, an ML model is composed of several artifacts that include the training dataset, ML algorithm, ML framework (e.g., PyTorch), model configuration (e.g., hyperparameters, weights), and metadata (e.g., license).
@@
The ML pipeline defines the sequence of operations or a workflow that transforms a model artifact at a particular stage of the ML lifecycle [48]. To support standardization and repeatability, the pipeline also facilitates workflow management and automation.
@@
In Atlas, metadata describes two mains aspects about an ML model. First, provenance metadata refers to the origin and history of custody of a model artifact, including its history of transformations as it traverses the ML lifecycle. Pipeline metadata, in turn, describes the ML systems and specifics about the operations that produced a model artifact.
An attestation in Atlas refers to any digitally signed metadata and serves as evidence for model artifact or ML system authenticity, integrity and provenance. Attestations may be generated by hardware or software.
@@ 3.2.1 Stakeholders
The Artifact Producers are individuals and organizations that create ML model artifacts to provide or sell them to other parties. Since the artifacts may represent intellectual property and/or make use of personally identifying information (PII), producers have business and regulatory reasons to preserve their confidentiality. To save costs, artifact producers often outsource the operation of ML systems to external service providers.
ML-as-a-Service (MLaaS) providers operate and maintain the compute infrastructure needed to run ML systems. MLaaS providers may offer ML-specific services that leverage general-purpose compute (e.g., [49, 50, 51]), or provide special-purpose systems (e.g., [37, 52]) that can build and run third-party ML systems.
A Hub is a system that stores and distributes model artifacts. Thus, model pipelines typically ingest and output artifacts to and from hubs during their execution, enabled by interfaces exposed by MLaaS providers. Hubs may be operated by artifact producers themselves or by third-party vendors, containing open or closed source artifacts (e.g., [53, 54]).
A Transparency Service in Atlas is responsible for generating, storing and distributing the metadata necessary to verify the authenticity, integrity and provenance of model artifacts. We envision model vendors and independent parties operating transparency services in practice.
Transparency services interface with MLaaS providers through attestation clients that run alongside ML systems to obtain and attest provenance and pipeline metadata. On the server side, a transparency log contains the known good values (i.e., golden values) of model artifacts and ML system components, submitted by model producers and MLaaS providers, as well as attestations collected by the clients throughout the ML lifecycle.
Verification Services evaluate or audit a particular ML model’s lifecycle with the goal of detecting unintended or malicious tampering with the model at any stage. In practice, model users, vendors or regulatory entities may operate verification services.
Using the golden values111 Golden values should be independently verified to establish their trustworthiness. Current approaches for auditing golden values include reproducibility [55] and endorsements [56]. Atlas is agnostic to the chosen method and assumes that evidence of golden value verification can be made available via a transparency service. and attestations obtained from a transparency service, a verification service evaluates each ML pipeline and artifact of interest against a set of expectations. For example, a model producer may check that the MLaaS provider ran the expected pipeline code, or a model user may verify that a fine-tuned model was produced from the expected foundation model.
@@ 3.2.2 ML Lifecycle
In Atlas, we consider four high-level stages in the ML lifecycle. Each builds upon the outputs and feedback from the others, forming a continuous cycle in which models evolve based on real-world usage.
1. Data processing: Raw data is collected, sanitized and processed into smaller units (e.g., tokens) and collated into a structure ingestable during training or evaluation.
2. Training: A training algorithm processes a given dataset using an ML system. The output is an ML model.
3. Evaluation: Following training, model properties like its performance and accuracy undergo further fine-tuning and evaluation using a testing dataset.
4. Deployment: After training and evaluation, an ML model is deployed to a production system configured for inferencing. New data obtained from clients during inference are sent back to a data processing pipeline to enhance the training dataset and the model. Model use must comply with local laws or corporate policies.
@@ 3.3 Threat Model
We consider an adversary whose goal is to produce a tampered artifact, e.g., containing a hidden malicious component, so that a transparency service generates a legitimate signature on the artifact or its metadata.
Thus, Atlas aims to detect such tampering introduced via the ML supply chain.222 Analogously to software correctness (which also applies to ML algorithms and systems), establishing dataset benignity, model quality and safety is a complementary area of research that relies on certifications, e.g., adversarial robustness [57], differential privacy [58], or poisoning [59]. We leave extending Atlas with such mechanisms as future work. Specifically, we consider tampering by MLaaS providers, hubs and artifact producers, while model users, transparency and verification services are trusted in Atlas.
Compromised MLaaS providers and hubs may involve malicious insiders, or external adversaries seeking to subvert these systems by exploiting vulnerable components. Given their central position in the lifecycle, MLaaS providers and hubs may thus be able to compromise model integrity at various stages.
For example, a malicious MLaaS provider can poison the training data during the curation step of the data processing stage leading to backdoors. A compromised hub may, for instance, present a dataset or model with a mismatched signature (e.g., to a different model, or any of its component artifacts) to a model user or MLaaS provider, so that pipelines in subsequent stages of the ML lifecycle may ingest compromised dependencies.
As a result, these compromises propagate through the ML lifecycle if they go undetected, ultimately leading to vulnerable ML models at the deployment stage. This risk is exacerbated if a hub colludes with an MLaaS provider to introduce or accept compromised ML pipeline inputs.
Artifact producers, on the other hand, may seek to compromise ML models to bypass regulatory requirements, introduce exploitable vulnerabilities or steal private information for profit (e.g, [60]). Thus, producers may collude with other untrusted stakeholders, or intentionally inaccurately declare their dependencies, to undermine the integrity of their artifacts.